NvidiaIntely (nvidiaintely):

Hannah is growing a garden in her backyard. It is in the shape of a square with side lengths of 7 feet. Her husband said that he can grow the same amount of crops in half the space. What will the side lengths of his garden measure? 7 feet 3.5 feet 7.5 feet 14 feet

NvidiaIntely (nvidiaintely):

@Kevin

OpenStudy (kevin):

the garden must be bigger than the square in it.. 7 * 2 = 14

OpenStudy (kevin):

wait

NvidiaIntely (nvidiaintely):

ok

NvidiaIntely (nvidiaintely):

The ratio of the areas of two similar polygons is 49:36. If the perimeter of the first polygon is 15 cm, what is the perimeter of the second polygon? Round to the nearest hundredth. 11.76 cm 11.02 cm 13.25 cm 12.85 cm

NvidiaIntely (nvidiaintely):

answer that right away?

NvidiaIntely (nvidiaintely):

ik I asked that already but lost it :P

OpenStudy (kevin):

perimeter of square = 7*4 = 28 The ratio of perimeter of square and the garden : 28 = 1/2* n n = 56 length side of the garden = 56/4 = 14

OpenStudy (kevin):

it's really 14

NvidiaIntely (nvidiaintely):

ok got that.